A solution prepared by adding 10g of an unknown non-volatile, non-electrolyte solute to 100 g of water. The Boiling point of the solution is elevated by 0.433 0C above the normal boiling point of the solvent. What is the molar mass of the unknown substance?

1
Expert's answer
2022-02-18T03:36:04-0500

∆Tb = i x Kb x m

∆Tb =Tb solution - Tb pure = change in boiling point

Kb = boiling point elevation constant

i = Vant Hoff factor of solute

m = molality = mol solute / kg solvent

∆Tb = 0.433 °C

Kb = 0.51 °C/m

For non-volatile, non-electrolyte solute i = 1

m = ∆Tb / (i x Kb) = 0.433 °C / (1 x 0.51 °C/m) = 0.849 mol/kg

m = 10 g / (Mr(solute) x 0.1 kg) = 0.849 mol/kg

Mr(solute) = 10 g / (0.849 mol/kg x 0.1 kg) = 117.8 g/mol